# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 5.1 Problem Formulation
where $\mu\in[0,1]$, $q:[0,1]\times\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{n}$ is a smooth path of the system states, $\varpi:\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{p}$ is a time-varying parameter, $\textsl{g}:\mathbb{R}^{n}\times\mathbb{R}^{p}\times\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{n}$ is a known smooth function which renders $\dot{q}=\textsl{g}(q,\varpi,t)$ contracting with respect to $q$, $d:[0,1]\times\mathbb{R}^{p}\times\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{n}$ with $\sup_{\mu,\varpi,t}\|\partial d/\partial\mu\|=\bar{d}\in\mathbb{R}_{\geq 0}$ is unknown external disturbance parameterized by $\mu$ as in Theorem 2.4, and $\Delta_{L}:\mathbb{R}^{p}\times\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{n}$ is the part to be learned with machine learning-based methodologies such as neural networks Hornik et al. [1989], Cybenko [1989], Funahashi [1989]. We can also formulate this in a stochastic setting as follows:
@@
where $\mathscr{W}:\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{w}$ is a Wiener process, and $G:[0,1]\times\mathbb{R}^{p}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times w}$ is a matrix-valued function parameterized by $\mu$ with $\sup_{\mu,\varpi,t}\|\partial G/\partial\mu\|_{F}=\bar{g}\in\mathbb{R}_{\geq 0}$ as in Theorem 2.5. The objective of learning is to find $\Delta_{L}$ which satisfies the following, assuming that $\varpi\in\mathcal{S}_{\varpi}\subseteq\mathbb{R}^{p}$ and $t\in\mathcal{S}_{t}\subseteq\mathbb{R}_{\geq 0}$ for some compact sets $\mathcal{S}_{\varpi}$ and $\mathcal{S}_{t}$:
@@
Examples 5.15.3 illustrate how the problem (5.3) with the nonlinear systems (5.1) and (5.2) can be used to describe several learning-based and data-driven control problems to be discussed in the subsequent sections, regarding $\varpi$ as $x$, $_{d}, x^\hat{x}$, $u_{d}$, etc.
Example 5.1 (Tracking control)
Let us consider the problem of learning a computationally-expensive (or unknown) feedback control law $u^{*}(x,x_{d},u_{d},t)$ that tracks a target trajectory $(x_{d},u_{d})$ given by $\dot{x}_{d}=f(x_{d},t)+B(x_{d},t)u_{d}$ as in (3.3), assuming that $f$ and $B$ are known. If the dynamics is perturbed by $d_{c}(x,t)$ as in (3.1), we have that
@@
where $u_{L}(x,x_{d},u_{d},t)$ denotes a learning-based control law that models $u^{*}(x,x_{d},u_{d},t)$. As long as $u^{*}$ renders the closed-loop dynamics $\dot{x}=f(x,t)+B(x,t)u^{*}(x,x_{d},u_{d},t)$ contracting with respect to $x$ Tsukamoto and Chung [2021], Manchester and Slotine [2017], Tsukamoto and Chung [2021], Singh et al. [2017], we can define the functions of (5.1) as follows:
@@
The learning objective here is to make $\|u_{L}(x,x_{d},u_{d},t)-u^{*}(x,x_{d},u_{d},t)\|$ as small as possible, which aligns with the aforementioned objective in (5.3). Note that if $\|B\|$ is bounded in the compact set $\mathcal{S}$ of (5.3), we can bound $\|\Delta_{L}\|$ of (5.6) with $\epsilon_{\ell 0}\neq 0$ and $\epsilon_{\ell 1}=0$ as to be explained in Remark 5.1. See Theorems 6.1, 6.3, and 7.1 for details.
Example 5.2 (State estimation)
Next, let us consider the problem of learning a computationally-expensive (or unknown) state estimator, approximating its estimation gain $L(\hat{x},t)$ by $L_{L}(\hat{x},t)$. If there is no disturbance in (4.3) and (4.4), we have that
@@
where $\dot{x}=f(x,t)$ is the true system, $y=h(x,t)$ is the measurement, and we assume that $f$ and $h$ are known. If $L$ is designed to render $\dot{q}=f(q,t)+L(\hat{x},t)(h(x,t)-h(q,t))$ contracting with respect to $q$ Tsukamoto and Chung [2021, 2021], Dani et al. [2015], Bonnabel and Slotine [2015], we could define the functions of (5.1) and (5.2) as follows:
@@
where $\varpi=[x^{\top},\hat{x}^{\top}]^{\top}$. It can be seen that (5.1) and (5.2) with the relations (5.9) and (5.10) indeed has $q(\mu=0,t)=x$ and $q(\mu=1,t)=\hat{x}$ as its particular solutions when perturbed by deterministic and stochastic disturbances as in (4.3) and (4.4), respectively. We can bound $\|\Delta_{L}\|$ of (5.10) in the compact set of (5.3) with $\epsilon_{\ell 0}=0$ and $\epsilon_{\ell 1}\neq 0$ if $h$ is Lipschitz, and with $\epsilon_{\ell 0}\neq 0$ and $\epsilon_{\ell 1}=0$ if $h$ is bounded in $\mathcal{S}$, using the techniques to be discussed in Remark 5.1. See Theorems 6.2 and 6.3 for details.
Example 5.3 (System identification)
We can use the problem (5.3) with the systems (5.1) and (5.2) also if $f_{\mathrm{true}}$ of the underlying dynamics $\dot{x}^{*}=f_{\mathrm{true}}(x^{*},t)$ is unknown and learned by $\dot{x}=f_{L}(x,t)$. Since we have
@@
where $\varpi=x^{*}$, as long as $\dot{x}=f_{L}(x,t)$ is contracting with respect to $x$ Singh et al. [2020], Boffi et al. [2020]. Since $\Delta_{L}$ of (5.12) is the learning error itself, $\|\Delta_{L}\|$ can be bounded in $\mathcal{S}$ using the techniques of Remark 5.1. See Theorems 9.1 and 9.2 for details.
@@
As seen in Examples 5.15.2, $\Delta_{L}$ of (5.3) is typically given by a learning error, $\phi(z)-\varphi(z)$, multiplied by a bounded or Lipschitz continuous function, where $z=(\varpi,t)\in\mathcal{S}$ for a compact set $\mathcal{S}$, and $\phi$ is the true computationally-expensive/unknown function to be learned by $\varphi$ as in (5.6), (5.10), (5.12), etc.
Let $\mathcal{T}$ be a set of training data $\{(z_{i},\phi(z_{i}))\}_{i=1}^{N}$ sampled in $\mathcal{S}$. For systems with the true function $\phi$ and its approximation $\varphi$ being Lipschitz (e.g., by spectral normalization Miyato et al. [2018] to be discussed in Definition 6.3 or by using contraction theory as in Revay et al. [2020, 2021], Manchester et al. [2021]), we can analytically find a bound for $\|\phi(z)-\varphi(z)\|$ as follows if target data samples $z$ are in $B(r)=\{z\in\mathcal{S}|\sup_{z^{\prime}\in\mathcal{T}}\|z-z^{\prime}\|\leq r\}$ for some $r\in\mathbb{R}_{\geq 0}$:
@@
where $L_{\phi}$ and $L_{\varphi}$ are the Lipschitz constants of $\phi$ and $\varphi$, respectively. The term $\sup_{z^{\prime}\in\mathcal{T}}\|\phi(z^{\prime})-\varphi(z^{\prime})\|$ can then be bounded by a constant, e.g., by using a deep robust regression model as proven in Chen et al. [2016], Liu et al. [2020] with spectral normalization, under standard training data distribution assumptions.
Deep Neural Networks (DNNs) have been shown to generalize well to the set of unseen events that are from almost the same distribution as their training set Zhang et al. [2017], He et al. [2016], Shi et al. [2019], Dziugaite and Roy [2017], Neyshabur et al. [2017], and consequently, obtaining a tighter and more general upper bound for the learning error as in (5.13) has been an active field of research Bartlett et al. [2017], Neyshabur et al. [2018], Liu et al. [2020], Chen et al. [2016]. Thus, the condition (5.3) has become a common assumption in analyzing the performance of learning-based and data-driven control techniques Shi et al. [2019], Taylor et al. [2019], Cheng et al. [2019], McKinnon and Schoellig [2019], Tsukamoto and Chung [2021], Tsukamoto et al. [2021, 2021], Shi et al. [2020, 2021].
5.2 Formal Stability Guarantees via Contraction Theory
@@
Suppose that $\Delta_{L}$ of (5.1) is learned to satisfy (5.3), and that $\textsl{g}(q,\varpi,t)$ of (5.1) is Lipschitz with respect to $q$ with its 2-norm Lipschitz constant $L_{\textsl{g}}\in\mathbb{R}_{\geq 0}$, i.e.,
@@
See the Gronwall-Bellman inequality [Khalil, 2002, p. 651] and Theorem 3.4 of [Khalil, 2002, pp. 96-97]. $\qed$
The bound obtained in Theorem 5.1 is useful in that it gives mathematical guarantees even for naive learning-based frameworks without a contracting property (e.g., it can be used to prove safety in the learning-based Model Predictive Control (MPC) framework Hewing et al. [2020]). However, the exponential term $e^{(L_{\textsl{g}}+\epsilon_{\ell 1})t}$ in (5.15) causes the upper bound to diverge, which could result in more conservative automatic control designs than necessary.
In contrast, contraction theory gives an upper bound on the trajectory tracking error $\|\mathtt{e}(t)\|$ which is exponentially bounded linearly in the learning error, even under the presence of external disturbances Tsukamoto and Chung [2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2021], Tsukamoto et al. [2021].
@@
Let us consider the virtual system of a smooth path $q(\mu,t)$ in (5.1) and suppose that $\Delta_{L}$ of (5.1) is learned to satisfy (5.3). If the condition $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32) holds and the system (5.1) with $\Delta_{L}=0$ and $d=0$ is contracting, i.e.,
@@
of Theorem 2.1 or 2.2 holds for $M\succ 0$ that defines a contraction metric with the contraction rate $\alpha$, and if the learning error $\epsilon_{\ell 1}$ of (5.3) is sufficiently small to satisfy
@@
where $\mathtt{e}(t)=\xi_{1}(t)-\xi_{0}(t)$ with $\xi_{0}(t)=q(0,t)$ and $\xi_{1}(t)=q(1,t)$ for $q$ of (5.1), $\bar{d}=\sup_{\mu,\varpi,t}\|\partial d/\partial\mu\|$ as in (5.1), $\mathcal{S}$ is the compact set of Theorem 5.1, $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$ are the learning errors of (5.3), and $V_{\ell}(t)=\int_{\xi_{0}}^{\xi_{1}}\|\Theta(q(t),t)\delta q(t)\|$ for $M=\Theta^{\top}\Theta$ as in Theorem 2.3.
@@
for all $\mu\in[0,1]$ and $(\varpi,t)\in\mathcal{S}$ due to $\sup_{\mu,\varpi,t}\|\partial d/\partial\mu\|=\bar{d}$, the direct application of Theorem 2.4 to the system (5.1), along with the condition (5.17), yields (5.18). $\qed$
Using Theorem 2.5, we can easily derive a stochastic counterpart of Theorem 5.2 for the system (5.2Tsukamoto et al. [2021, 2021].
@@
Consider the virtual system of a smooth path $q(\mu,t)$ in (5.2) and suppose $\Delta_{L}$ of (5.2) is learned to satisfy (5.3). If $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32) holds and the system (5.2) is contracting, i.e.,
@@
of Theorem 2.5 holds, and if the learning error $\epsilon_{\ell 1}$ of (5.3) and an arbitrary constant $\alpha_{d}\in\mathbb{R}_{>0}$ (see (5.24)) are selected to satisfy
@@
where $\mathtt{e}(t)=\xi_{1}(t)-\xi_{0}(t)$ with $\xi_{0}(t)=q(0,t)$ and $\xi_{1}(t)=q(1,t)$ for $q$ of (5.2), $\mathcal{S}$ is the compact set given in Theorem 5.1, $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$ are the learning errors of (5.3), $V_{s\ell}(t)=\int_{\xi_{0}}^{\xi_{1}}\|\Theta(q(t),t)\delta q(t)\|^{2}$ for $M=\Theta^{\top}\Theta$ as given in Theorem 2.3, and $C=\bar{g}^{2}({2}{\alpha_{G}}^{-1}+1)+\epsilon_{\ell 0}^{2}\alpha_{d}^{-1}$ for $\bar{g}=\sup_{\mu,\varpi,t}\|\partial G/\partial\mu\|_{F}$ of (5.2) with an arbitrary constant $\alpha_{G}\in\mathbb{R}_{>0}$ as in Theorem 2.5. Furthermore, the probability that $\|\mathtt{e}\|$ is greater than or equal to $\varepsilon\in\mathbb{R}_{>0}$ is given as
@@
Computing $\mathscr{L}V_{s\ell}$ with the virtual system (5.2) as in the proof of Theorem 2.5, we get an additional term $2\int_{0}^{1}\partial_{\mu}q^{\top}M\Delta_{L}$. Using the learning error assumption (5.3) to have $\|\Delta_{L}\|\leq\epsilon_{\ell 0}+(\epsilon_{\ell 1}/\sqrt{\underline{m}})\int_{0}^{1}\|\Theta\partial_{\mu}q\|d\mu$ as in (5.19), we have that
@@
where $V_{\ell}=\int_{0}^{1}\|\Theta\partial_{\mu}q\|d\mu$ as in Theorem 2.3, and the relation $2ab\leq\alpha_{d}^{-1}a^{2}+\alpha_{d}b^{2}$, which holds for any $a,b\in\mathbb{R}$ and $\alpha_{d}\in\mathbb{R}_{>0}$ ($a=\epsilon_{\ell 0}\sqrt{\overline{m}}$ and $b=V_{\ell}$ in this case), is used to obtain the second inequality. Since we have $V_{\ell}^{2}\leq V_{s\ell}$ as proven in Theorem 2.3, selecting $\alpha_{d}$ and $\epsilon_{\ell 1}$ sufficiently small to satisfy (5.21) gives the desired relations (5.22) and (5.23) due to Theorem 2.5. $\qed$
@@
As discussed in Theorems 5.2 and 5.3, using contraction theory for learning-based and data-driven control, we can formally guarantee the system trajectories to stay in a tube with an exponentially convergent bounded radius centered around the target trajectory, even with the external disturbances and learning errors $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$. The exponential bounds (5.18), (5.22), and (5.23) become tighter as we achieve smaller $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$ using more training data for verifying (5.3) (see Remark 5.1). It is also worth noting that the steady-state bounds of (5.18) and (5.22) are again some functions of the condition number $\chi=\overline{m}/\underline{m}$ as in Theorems 4.1 and 4.4, which renders the aforementioned CV-STEM approach valid and effective also in the learning-based frameworks. Theorems 5.2 and 5.3 play a central role in providing incremental exponential stability guarantees for learning-based and data-driven automatic control to be introduced in Sec. 68.
5.3 Motivation for Neural Contraction Metrics

### v8

@@ 5.1. Problem Formulation
where $\mu\in[0,1]$, $q:[0,1]\times\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{n}$ is a smooth path of the system states, $\varpi:\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{p}$ is a time-varying parameter, $\textsl{g}:\mathbb{R}^{n}\times\mathbb{R}^{p}\times\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{n}$ is a known smooth function which renders $\dot{q}=\textsl{g}(q,\varpi,t)$ contracting with respect to $q$, $d:[0,1]\times\mathbb{R}^{p}\times\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{n}$ with $\sup_{\mu,\varpi,t}\|\partial d/\partial\mu\|=\bar{d}\in\mathbb{R}_{\geq 0}$ is unknown external disturbance parameterized by $\mu$ as in Theorem LABEL:THM:Thm:Robust_contraction_original, and $\Delta_{L}:\mathbb{R}^{p}\times\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{n}$ is the part to be learned with machine learning-based methodologies such as neural networks neural1, neural2, neural3. We can also formulate this in a stochastic setting as follows:
@@
where $\mathscr{W}:\mathbb{R}_{\geq 0}\mapsto\mathbb{R}^{w}$ is a Wiener process, and $G:[0,1]\times\mathbb{R}^{p}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n\times w}$ is a matrix-valued function parameterized by $\mu$ with $\sup_{\mu,\varpi,t}\|\partial G/\partial\mu\|_{F}=\bar{g}\in\mathbb{R}_{\geq 0}$ as in Theorem LABEL:THM:Thm:robuststochastic. The objective of learning is to find $\Delta_{L}$ which satisfies the following, assuming that $\varpi\in\mathcal{S}_{\varpi}\subseteq\mathbb{R}^{p}$ and $t\in\mathcal{S}_{t}\subseteq\mathbb{R}_{\geq 0}$ for some compact sets $\mathcal{S}_{\varpi}$ and $\mathcal{S}_{t}$:
@@
Examples LABEL:EX:ex:learning_prob1LABEL:EX:ex:learning_prob3 illustrate how the problem (5.3) with the nonlinear systems (5.1) and (5.2) can be used to describe several learning-based and data-driven control problems to be discussed in the subsequent sections, regarding $\varpi$ as $x$, $_{d}, x^\hat{x}$, $u_{d}$, etc.
Example 5.1
Learning-based Tracking Controlex:learning_prob1 Let us consider the problem of learning a computationally-expensive (or unknown) feedback control law $u^{*}(x,x_{d},u_{d},t)$ that tracks a target trajectory $(x_{d},u_{d})$ given by $\dot{x}_{d}=f(x_{d},t)+B(x_{d},t)u_{d}$ as in (3.3), assuming that $f$ and $B$ are known. If the dynamics is perturbed by $d_{c}(x,t)$ as in (3.1), we have that
@@
where $u_{L}(x,x_{d},u_{d},t)$ denotes a learning-based control law that models $u^{*}(x,x_{d},u_{d},t)$. As long as $u^{*}$ renders the closed-loop dynamics $\dot{x}=f(x,t)+B(x,t)u^{*}(x,x_{d},u_{d},t)$ contracting with respect to $x$ mypaperTAC, ccm, ncm, 7989693, we can define the functions of (5.1) as follows:
@@
The learning objective here is to make $\|u_{L}(x,x_{d},u_{d},t)-u^{*}(x,x_{d},u_{d},t)\|$ as small as possible, which aligns with the aforementioned objective in (5.3). Note that if $\|B\|$ is bounded in the compact set $\mathcal{S}$ of (5.3), we can bound $\|\Delta_{L}\|$ of (5.6) with $\epsilon_{\ell 0}\neq 0$ and $\epsilon_{\ell 1}=0$ as to be explained in Remark 5.1. See Theorems LABEL:THM:Thm:NCMstability1, LABEL:THM:Thm:NCMstability2, and LABEL:THM:Thm:lagros_stability for details.
Example 5.2
Learning-based State Estimationex:learning_prob2 Next, let us consider the problem of learning a computationally-expensive (or unknown) state estimator, approximating its estimation gain $L(\hat{x},t)$ by $L_{L}(\hat{x},t)$. If there is no disturbance in (4.3) and (4.4), we have that
@@
where $\dot{x}=f(x,t)$ is the true system, $y=h(x,t)$ is the measurement, and we assume that $f$ and $h$ are known. If $L$ is designed to render $\dot{q}=f(q,t)+L(\hat{x},t)(h(x,t)-h(q,t))$ contracting with respect to $q$ ncm, mypaperTAC, Ref:Stochastic, 6849943, we could define the functions of (5.1) and (5.2) as follows:
@@
where $\varpi=[x^{\top},\hat{x}^{\top}]^{\top}$. It can be seen that (5.1) and (5.2) with the relations (5.9) and (5.10) indeed has $q(\mu=0,t)=x$ and $q(\mu=1,t)=\hat{x}$ as its particular solutions when perturbed by deterministic and stochastic disturbances as in (4.3) and (4.4), respectively. We can bound $\|\Delta_{L}\|$ of (5.10) in the compact set of (5.3) with $\epsilon_{\ell 0}=0$ and $\epsilon_{\ell 1}\neq 0$ if $h$ is Lipschitz, and with $\epsilon_{\ell 0}\neq 0$ and $\epsilon_{\ell 1}=0$ if $h$ is bounded in $\mathcal{S}$, using the techniques to be discussed in Remark 5.1. See Theorems LABEL:THM:Thm:NCMstability1_est and LABEL:THM:Thm:NCMstability2 for details.
Example 5.3
Learning-based System Identificationex:learning_prob3 We can use the problem (5.3) with the systems (5.1) and (5.2) also if $f_{\mathrm{true}}$ of the underlying dynamics $\dot{x}^{*}=f_{\mathrm{true}}(x^{*},t)$ is unknown and learned by $\dot{x}=f_{L}(x,t)$. Since we have
@@
where $\varpi=x^{*}$, as long as $\dot{x}=f_{L}(x,t)$ is contracting with respect to $x$ 47710, boffi2020learning. Since $\Delta_{L}$ of (5.12) is the learning error itself, $\|\Delta_{L}\|$ can be bounded in $\mathcal{S}$ using the techniques of Remark 5.1. See Theorems LABEL:THM:Thm:NCMstability_modelfree and LABEL:THM:Thm:neurallander for details.
@@
As seen in Examples LABEL:EX:ex:learning_prob1LABEL:EX:ex:learning_prob2, $\Delta_{L}$ of (5.3) is typically given by a learning error, $\phi(z)-\varphi(z)$, multiplied by a bounded or Lipschitz continuous function, where $z=(\varpi,t)\in\mathcal{S}$ for a compact set $\mathcal{S}$, and $\phi$ is the true computationally-expensive/unknown function to be learned by $\varphi$ as in (5.6), (5.10), (5.12), etc.
Let $\mathcal{T}$ be a set of training data $\{(z_{i},\phi(z_{i}))\}_{i=1}^{N}$ sampled in $\mathcal{S}$. For systems with the true function $\phi$ and its approximation $\varphi$ being Lipschitz (e.g., by spectral normalization miyato2018spectral to be discussed in Definition LABEL:DEF:Def:SN or by using contraction theory as in revay2020lipschitz, revay2021recurrent, cdc_systemid), we can analytically find a bound for $\|\phi(z)-\varphi(z)\|$ as follows if target data samples $z$ are in $B(r)=\{z\in\mathcal{S}|\sup_{z^{\prime}\in\mathcal{T}}\|z-z^{\prime}\|\leq r\}$ for some $r\in\mathbb{R}_{\geq 0}$:
@@
where $L_{\phi}$ and $L_{\varphi}$ are the Lipschitz constants of $\phi$ and $\varphi$, respectively. The term $\sup_{z^{\prime}\in\mathcal{T}}\|\phi(z^{\prime})-\varphi(z^{\prime})\|$ can then be bounded by a constant, e.g., by using a deep robust regression model as proven in pmlr-v51-chen16d, nnbound with spectral normalization, under standard training data distribution assumptions.
Deep Neural Networks (DNNs) have been shown to generalize well to the set of unseen events that are from almost the same distribution as their training set rethinknet, dnnlearn, neurallander, DBLP:conf/uai/DziugaiteR17, NIPS2017_10ce03a1, and consequently, obtaining a tighter and more general upper bound for the learning error as in (5.13) has been an active field of research marginbounds1, marginbounds2, nnbound, pmlr-v51-chen16d. Thus, the condition (5.3) has become a common assumption in analyzing the performance of learning-based and data-driven control techniques neurallander, 8967820, pmlr-v97-cheng19a, 8651519, lagros, ancm, cdc_ncm, swarm1, swarm2.
5.2. Formal Stability Guarantees via Contraction Theory
@@
Naive Learning-based ControlThm:naive_learning Suppose that $\Delta_{L}$ of (5.1) is learned to satisfy (5.3), and that $\textsl{g}(q,\varpi,t)$ of (5.1) is Lipschitz with respect to $q$ with its 2-norm Lipschitz constant $L_{\textsl{g}}\in\mathbb{R}_{\geq 0}$, i.e.,
@@
See the Gronwall-Bellman lemma [Khalil:1173048, p. 651] and Theorem 3.4 of [Khalil:1173048, pp. 96-97]. $\qed$
The bound obtained in Theorem LABEL:THM:Thm:naive_learning is useful in that it gives mathematical guarantees even for naive learning-based frameworks without a contracting property (e.g., it can be used to prove safety in the learning-based Model Predictive Control (MPC) framework learningmpc). However, the exponential term $e^{(L_{\textsl{g}}+\epsilon_{\ell 1})t}$ in (5.15) causes the upper bound to diverge, which could result in more conservative automatic control designs than necessary.
In contrast, contraction theory gives an upper bound on the trajectory tracking error $\|\mathtt{e}(t)\|$ which is exponentially bounded linearly in the learning error, even under the presence of external disturbances ncm, ancm, lagros, cdc_ncm.
@@
Learning-based Control with Deterministic ContractionThm:contraction_learning Let us consider the virtual system of a smooth path $q(\mu,t)$ in (5.1) and suppose that $\Delta_{L}$ of (5.1) is learned to satisfy (5.3). If the condition $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32) holds and the system (5.1) with $\Delta_{L}=0$ and $d=0$ is contracting, i.e.,
@@
of Theorem LABEL:THM:Thm:contraction or LABEL:THM:Thm:partial_contraction holds for $M\succ 0$ that defines a contraction metric with the contraction rate $\alpha$, and if the learning error $\epsilon_{\ell 1}$ of (5.3) is sufficiently small to satisfy
@@
where $\mathtt{e}(t)=\xi_{1}(t)-\xi_{0}(t)$ with $\xi_{0}(t)=q(0,t)$ and $\xi_{1}(t)=q(1,t)$ for $q$ of (5.1), $\bar{d}=\sup_{\mu,\varpi,t}\|\partial d/\partial\mu\|$ as in (5.1), $\mathcal{S}$ is the compact set of Theorem LABEL:THM:Thm:naive_learning, $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$ are the learning errors of (5.3), and $V_{\ell}(t)=\int_{\xi_{0}}^{\xi_{1}}\|\Theta(q(t),t)\delta q(t)\|$ for $M=\Theta^{\top}\Theta$ as in Theorem LABEL:THM:Thm:path_integral.
@@
for all $\mu\in[0,1]$ and $(\varpi,t)\in\mathcal{S}$ due to $\sup_{\mu,\varpi,t}\|\partial d/\partial\mu\|=\bar{d}$, the direct application of Theorem LABEL:THM:Thm:Robust_contraction_original to the system (5.1), along with the condition (5.17), yields (5.18). $\qed$
Using Theorem LABEL:THM:Thm:robuststochastic, we can easily derive a stochastic counterpart of Theorem LABEL:THM:Thm:contraction_learning for the system (5.2nscm, cdc_ncm.
@@
Learning-based Control with Stochastic ContractionThm:contraction_learning_sto Consider the virtual system of a smooth path $q(\mu,t)$ in (5.2) and suppose $\Delta_{L}$ of (5.2) is learned to satisfy (5.3). If $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32) holds and the system (5.2) is contracting, i.e.,
@@
of Theorem LABEL:THM:Thm:robuststochastic holds, and if the learning error $\epsilon_{\ell 1}$ of (5.3) and an arbitrary constant $\alpha_{d}\in\mathbb{R}_{>0}$ (see (5.24)) are selected to satisfy
@@
where $\mathtt{e}(t)=\xi_{1}(t)-\xi_{0}(t)$ with $\xi_{0}(t)=q(0,t)$ and $\xi_{1}(t)=q(1,t)$ for $q$ of (5.2), $\mathcal{S}$ is the compact set given in Theorem LABEL:THM:Thm:naive_learning, $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$ are the learning errors of (5.3), $V_{s\ell}(t)=\int_{\xi_{0}}^{\xi_{1}}\|\Theta(q(t),t)\delta q(t)\|^{2}$ for $M=\Theta^{\top}\Theta$ as given in Theorem LABEL:THM:Thm:path_integral, and $C=\bar{g}^{2}({2}{\alpha_{G}}^{-1}+1)+\epsilon_{\ell 0}^{2}\alpha_{d}^{-1}$ for $\bar{g}=\sup_{\mu,\varpi,t}\|\partial G/\partial\mu\|_{F}$ of (5.2) with an arbitrary constant $\alpha_{G}\in\mathbb{R}_{>0}$ as in Theorem LABEL:THM:Thm:robuststochastic. Furthermore, the probability that $\|\mathtt{e}\|$ is greater than or equal to $\varepsilon\in\mathbb{R}_{>0}$ is given as
@@
Computing $\mathscr{L}V_{s\ell}$ with the virtual system (5.2) as in the proof of Theorem LABEL:THM:Thm:robuststochastic, we get an additional term $2\int_{0}^{1}\partial_{\mu}q^{\top}M\Delta_{L}$. Using the learning error assumption (5.3) to have $\|\Delta_{L}\|\leq\epsilon_{\ell 0}+(\epsilon_{\ell 1}/\sqrt{\underline{m}})\int_{0}^{1}\|\Theta\partial_{\mu}q\|d\mu$ as in (5.19), we have that
@@
where $V_{\ell}=\int_{0}^{1}\|\Theta\partial_{\mu}q\|d\mu$ as in Theorem LABEL:THM:Thm:path_integral, and the relation $2ab\leq\alpha_{d}^{-1}a^{2}+\alpha_{d}b^{2}$, which holds for any $a,b\in\mathbb{R}$ and $\alpha_{d}\in\mathbb{R}_{>0}$ ($a=\epsilon_{\ell 0}\sqrt{\overline{m}}$ and $b=V_{\ell}$ in this case), is used to obtain the second inequality. Since we have $V_{\ell}^{2}\leq V_{s\ell}$ as proven in Theorem LABEL:THM:Thm:path_integral, selecting $\alpha_{d}$ and $\epsilon_{\ell 1}$ sufficiently small to satisfy (5.21) gives the desired relations (5.22) and (5.23) due to Theorem LABEL:THM:Thm:robuststochastic. $\qed$
@@
As discussed in Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto, using contraction theory for learning-based and data-driven control, we can formally guarantee the system trajectories to stay in a tube with an exponentially convergent bounded radius centered around the target trajectory, even with the external disturbances and learning errors $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$. The exponential bounds (5.18), (5.22), and (5.23) become tighter as we achieve smaller $\epsilon_{\ell 0}$ and $\epsilon_{\ell 1}$ using more training data for verifying (5.3) (see Remark 5.1). It is also worth noting that the steady-state bounds of (5.18) and (5.22) are again some functions of the condition number $\chi=\overline{m}/\underline{m}$ as in Theorems LABEL:THM:Thm:CVSTEMobjective and LABEL:THM:Thm:CV-STEMobjective-est, which renders the aforementioned CV-STEM approach valid and effective also in the learning-based frameworks. Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto play a central role in providing incremental exponential stability guarantees for learning-based and data-driven automatic control to be introduced in Sec. 68.
5.3. Motivation for Neural Contraction Metrics